HUBBLE SPACE TELESCOPE DAILY REPORT #3098 PERIOD COVERED: 0000Z (UTC) 04/18/02 - 0000Z (UTC) 04/19/02 Daily Status Report as of 109/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed Five Sets ACS/WFC 9018 (Flat Field Stability) The Advanced Camera for Surveys (WFC) was used to assess the stability and uniformity of the low-frequency flat fields {L-flat} of the WFC detector by using multiple pointing observations of the globular clusters 47 Tucanae {NGC104}, thus imaging moderately dense stellar fields. By placing the same star over different portions of the detectors and measuring relative changes in its brightness it will be possible to determine local variations in the response of the detectors. There were no reported problems. 1.2 Completed Two Sets of STIS/CCD 8902 (Dark Monitor-Part 2) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the darks. No problems were reported. 1.3 Completed Three Sets of WF/PC-2 8937 (Cycle 9 Supplemental Darks pt2/3) The WF/PC-2 was used obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. No problems were encountered. 1.4 Completed Five Sets of WF/PC-2 8950 (SM3B Cool Down, Contamination Monitor, and Focus Check) The WF/PC-2 was used to monitor the planned cooldown to -88C, followed by UV monitors, focus checks, as well as decons and associated monitors. No problems were noted. 1.5 Completed STIS/MA2 9573 (NUV-MAMA Daily Dark Monitor) The Space Telescope Imaging Spectrograph (MA2) was used to perform daily monitoring of the NUV MAMA detector dark noise in order to monitor the effects of thermal changes on the NUV dark rate. No problems occurred. 1.6 Completed Three Sets of ACS/WFC/HRC 8947 (Weekly Test) The Advanced Camera for Surveys (WFC and HRC) was used to perform basic tests to monitor, the read noise, the development of hot pixels and test for any source of noise in ACS CCD detectors. This program will be executed at least once a day for the entire lifetime of ACS. There were no problems reported. 1.7 Completed STIS/CCD 8906 (Hot Pixel Annealing) The Space Telescope Imaging Spectrograph (CCD) was used to measure the effectiveness of the CCD hot pixel annealing process by measuring the dark current behavior before and after annealing and by searching for any window contamination effects. In addition, CTE performance is examined by looking for traps in a low signal level flat. There were no reported problems. 1.8 Completed STIS/CCD 8904 (Bias Monitor-Part 2) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4 in order to build up high-S/N superbiases and track the evolution of hot columns. The proposal completed with no anomalous activity. ************************************************** The following scheduled proposals did not execute due to the BALL decision to delay the ACS high voltage on commanding: ACS/SBC 9008 (MAMA Initial Turn-on and Anomalous Recovery Procedure) ACS/SBC 9009 (MAMA Fold Analysis) ************************************************** 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 7 Successful: 7 Scheduled Re-acquisitions: 13 Successful: 13 2.2 FHST Updates: Scheduled: 7 Successful: 7 2.3 Operations Notes: Using ROP IC-2, SSA transmitter #2 was turned on for the following intervals: 108/0906Z to 108/0926Z, 108/0949Z to 108/1013Z, and 108/1452Z to 108/1503Z. The 486 engineering status buffer limit for SESBSLD was modified at 108/0959Z as directed by ROP DF-18A. HSTAR 8616 was written when the AX Link 2 Temperature (EAXLK2T) flagged out-of-limits for one telemetry sample at 108/0635Z. The value was 9.19844. A TTR was written for a required re-transmit at 108/1325Z during a NSSC-1 load. Accordingly, per ROP NS-5, SI C&DH errors were reset at 108/1332Z. An operations request was used at 108/1459Z for tape recorder #2 management in preparation for battery #3 reconditioning activities. The SSR-1 EDAC error counter was cleared at 108/2259Z, using ROP SR-1A. -E SPA was unmasked from the TRSWCC stack at 109/0050Z per an operations request. Battery #3 reconditioning was completed at 109/0101Z. The Observer Rate Bias - Estimate 3 (QDVTHJ2) flagged out-of-limits at 109/053526Z with a value of 7.101e-5. The parameter was back in bounds at the conclusion of the following ZOE. ROP SR-9A was utilized to clear the SSR-3 EDAC error counter at 109/0355Z. The NICMOS instrument was commanded to BOOT, via SMS commanding, at 109/0708Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of Servicing Mission Orbital Verification and the gradual resumption of normal science observations and calibrations. /CAW